# Dependency Parsing is More Parameter-Efficient with Normalization

### v1

@@ 6 Conclusions
Departing from a state-of-the-art architecture for semantic dependency parsing, we were able to improve its performance on both semantic and syntactic dependency parsing on six datasets. On ERFGC, a dataset of directed acyclic semantic dependency graphs compiled from culinary recipes, our approach allowed us to beat the state-of-the-art performance achieved by Bhatt et al. bhatt_end–end_2024 . Moreover, our results showed that a single BiLSTM layer can be sufficient to match or surpass the results of state-of-the-art architectures with a decrease in trained parameters of up to 85%. In the case of SciERC, a challenging dataset for semantic dependency parsing, we found that the performance boost was particularly great when only training the biaffine scorer, without any BiLSTM layers. Moreover, for this challenging dataset, we find that scaling the predictions of the biaffine scorer can accelerate convergence speed even when fully fine-tuning models in the $100$ to $400M$ parameter range. In addition, for three of the datasets we also observed that the performance obtained with normalized and raw scores converged smoothly as the number of trained layers increased. This supported our claim that stacking BiLSTM layers mainly serves the purpose of producing an implicit regularization, and that this effect can be obtained by normalizing the scores, without any extra parameters.
@@
Due to the marginal performance increase reported for GNN-based dependency parsers ji_graph-based_2019 , iterative adjacency refinement methods in DP might suffer from limitations akin to the lack of normalization which we will explore in future work. This will also help us verify the actual effectiveness of triaffine two-hop and $k$-hop methods in general. Furthermore, we plan to extend our work to large scale graph inference tasks which have been limited by the lack of parameter efficient methods, such as long form discourse parsing tasks. With regard to score scaling, we also wish to verify whether the positive effects of this normalization can carry over to other tasks. Finally, further explorations in model efficiency are warranted, given our findings, e.g., via pruning of model parameters.
Appendix A Full Results
Table 4 reports the results for the best hyperparameter combinations for all three tasks of predicting tags, edges, and relations of the semantic dependency graphs. Note that the first part of Table 4 is identical to the bottom of Table 3. Since we already discussed the performance for labeled edges in Section 5, we forgo discussing them again in this appendix.
As regards unlabeled edges, using score normalization improves mean performance for all datasets, similarly to relations. This is expected, as unlabeled edges directly influence the prediction of the edge labels, together with entity class prediction. As we have already observed for relations, the performance boost provided by normalization is particularly great for SciERC at $L_{\psi}=0$.
It is also interesting to notice that for CoNLL04, the performance for the unlabeled edge prediction task is only slightly higher than that of the relations. In fact, as regards ADE, the opposite is true, with the labeled performance being seemingly higher. Due to the high variances, it would therefore seem that for these two datasets there is essentially no difference in difficulty between the labeled and unlabeled edge prediction tasks. In the case of ADE, this makes perfect sense, since there is only one possible relation. For CoNLL04, the amount of relation types is rather limited as well; therefore, it is sensible for the performance to also be similar in this case. Indeed, when looking at SciERC and ERFGC, the performance gap between labeled and unlabeled tasks is considerable. This is a strong indication that CoNLL04 is thus found somewhere in the middle, where the number of possible relations only slightly affects labeled performance.
As far as enEWT and SciDTB are concerned, the performance of edges (UAS) and relations (LAS) is also rather similar. Since in this case the predictions involve syntactic rather than semantic relations, the similar performance hints at relations being easier to predict in SynDP than in SemDP.
In the tagging task, very high standard deviation can be observed. In the case of ADE and CoNLL04, this could be caused by our choice of $\lambda_{1}=0.1$ being too low a value. Regarding SciERC, as already mentioned, most of the entities in the validation and test set are not found in the training set. This means it is not possible to train the model to recognize them, which results in the abysmal tagging performance reported in Table 4. This makes it challenging to train good tag embeddings which can help inform the edge and relation prediction tasks. Surprisingly, we notice that for CoNLL04 and ERFGC the tagging performance is also seemingly higher with score normalization. However, the overlap of the standard deviations is too high to be able to make any claims on the matter.
Appendix B Ablations
@@
Table 5: Tagger ablation results: performance for the best hyperparameters $(h_{\psi},d_{\text{MLP}})$. Best in bold, second-best underlined.
$a$ $L_{\psi}$ $\phi$ $\textbf{e}_{i}^{tag}$ ADE CoNLL04 SciERC ERFGC Mean
@@
$1$ 0 $\blacksquare$ $\blacksquare$ 0.541 ±$\pm 0.021$ 0.399 ±$\pm 0.024$ 0.147 ±$\pm 0.049$ 0.548 ±$\pm 0.010$ 0.515
1 $\blacksquare$ $\blacksquare$ 0.657 ±$\pm 0.011$ 0.556 ±$\pm 0.021$ 0.282 ±$\pm 0.009$ 0.676 ±$\pm 0.010$
2 $\blacksquare$ $\blacksquare$ 0.667 ±$\pm 0.011$ 0.573 ±$\pm 0.025$ 0.273 ±$\pm 0.010$ 0.694 ±$\pm 0.010$
3 $\blacksquare$ $\blacksquare$ 0.662 ±$\pm 0.027$ 0.562 ±$\pm 0.021$ 0.299 ±$\pm 0.023$ 0.705 ±$\pm 0.011$
$\frac{1}{\sqrt{d}}$ 0 $\blacksquare$ $\blacksquare$ 0.567 ±$\pm 0.014$ 0.438 ±$\pm 0.033$ 0.181 ±$\pm 0.027$ 0.612 ±$\pm 0.008$ 0.541
1 $\blacksquare$ $\blacksquare$ 0.668 ±$\pm 0.017$ 0.597 ±$\pm 0.015$ 0.299 ±$\pm 0.019$ 0.692 ±$\pm 0.009$
2 $\blacksquare$ $\blacksquare$ 0.676 ±$\pm 0.019$ 0.596 ±$\pm 0.014$ 0.312 ±$\pm 0.011$ 0.699 ±$\pm 0.009$
3 $\blacksquare$ $\blacksquare$ 0.686 ±$\pm 0.025$ 0.602 ±$\pm 0.017$ 0.320 ±$\pm 0.013$ 0.708 ±$\pm 0.008$
$1$ 0 $\blacksquare$ $\square$ 0.543 ±$\pm 0.013$ 0.402 ±$\pm 0.023$ 0.162 ±$\pm 0.019$ 0.554 ±$\pm 0.008$ 0.517
1 $\blacksquare$ $\square$ 0.674 ±$\pm 0.011$ 0.527 ±$\pm 0.026$ 0.275 ±$\pm 0.013$ 0.677 ±$\pm 0.005$
2 $\blacksquare$ $\square$ 0.672 ±$\pm 0.019$ 0.575 ±$\pm 0.009$ 0.293 ±$\pm 0.012$ 0.689 ±$\pm 0.011$
3 $\blacksquare$ $\square$ 0.657 ±$\pm 0.027$ 0.583 ±$\pm 0.011$ 0.301 ±$\pm 0.012$ 0.697 ±$\pm 0.007$
$\frac{1}{\sqrt{d}}$ 0 $\blacksquare$ $\square$ 0.563 ±$\pm 0.013$ 0.443 ±$\pm 0.019$ 0.188 ±$\pm 0.006$ 0.612 ±$\pm 0.003$ 0.534
1 $\blacksquare$ $\square$ 0.653 ±$\pm 0.023$ 0.565 ±$\pm 0.027$ 0.299 ±$\pm 0.020$ 0.687 ±$\pm 0.006$
2 $\blacksquare$ $\square$ 0.672 ±$\pm 0.017$ 0.592 ±$\pm 0.020$ 0.307 ±$\pm 0.008$ 0.702 ±$\pm 0.005$
3 $\blacksquare$ $\square$ 0.661 ±$\pm 0.022$ 0.593 ±$\pm 0.017$ 0.305 ±$\pm 0.014$ 0.708 ±$\pm 0.007$
$1$ 0 $\square$ $\blacksquare$ 0.550 ±$\pm 0.026$ 0.387 ±$\pm 0.030$ 0.157 ±$\pm 0.012$ 0.553 ±$\pm 0.006$ 0.514
1 $\square$ $\blacksquare$ 0.667 ±$\pm 0.022$ 0.532 ±$\pm 0.036$ 0.273 ±$\pm 0.013$ 0.673 ±$\pm 0.006$
2 $\square$ $\blacksquare$ 0.665 ±$\pm 0.021$ 0.565 ±$\pm 0.024$ 0.278 ±$\pm 0.027$ 0.694 ±$\pm 0.013$
3 $\square$ $\blacksquare$ 0.676 ±$\pm 0.021$ 0.558 ±$\pm 0.051$ 0.288 ±$\pm 0.012$ 0.706 ±$\pm 0.006$
$\frac{1}{\sqrt{d}}$ 0 $\square$ $\blacksquare$ 0.578 ±$\pm 0.022$ 0.437 ±$\pm 0.030$ 0.187 ±$\pm 0.014$ 0.611 ±$\pm 0.008$ 0.534
1 $\square$ $\blacksquare$ 0.651 ±$\pm 0.032$ 0.559 ±$\pm 0.029$ 0.301 ±$\pm 0.007$ 0.684 ±$\pm 0.003$
2 $\square$ $\blacksquare$ 0.673 ±$\pm 0.029$ 0.582 ±$\pm 0.017$ 0.310 ±$\pm 0.014$ 0.701 ±$\pm 0.008$
3 $\square$ $\blacksquare$ 0.659 ±$\pm 0.019$ 0.586 ±$\pm 0.026$ 0.324 ±$\pm 0.019$ 0.708 ±$\pm 0.012$
$1$ 0 $\square$ $\square$ 0.547 ±$\pm 0.019$ 0.402 ±$\pm 0.033$ 0.156 ±$\pm 0.029$ 0.549 ±$\pm 0.015$ 0.516
1 $\square$ $\square$ 0.651 ±$\pm 0.017$ 0.552 ±$\pm 0.012$ 0.288 ±$\pm 0.008$ 0.680 ±$\pm 0.007$
2 $\square$ $\square$ 0.664 ±$\pm 0.031$ 0.566 ±$\pm 0.012$ 0.288 ±$\pm 0.017$ 0.693 ±$\pm 0.011$
3 $\square$ $\square$ 0.663 ±$\pm 0.008$ 0.561 ±$\pm 0.014$ 0.291 ±$\pm 0.012$ 0.703 ±$\pm 0.007$
$\frac{1}{\sqrt{d}}$ 0 $\square$ $\square$ 0.566 ±$\pm 0.014$ 0.431 ±$\pm 0.015$ 0.182 ±$\pm 0.023$ 0.606 ±$\pm 0.012$ 0.534
1 $\square$ $\square$ 0.655 ±$\pm 0.015$ 0.567 ±$\pm 0.010$ 0.286 ±$\pm 0.018$ 0.678 ±$\pm 0.013$
2 $\square$ $\square$ 0.678 ±$\pm 0.014$ 0.591 ±$\pm 0.026$ 0.307 ±$\pm 0.007$ 0.702 ±$\pm 0.007$
3 $\square$ $\square$ 0.684 ±$\pm 0.022$ 0.595 ±$\pm 0.017$ 0.321 ±$\pm 0.016$ 0.698 ±$\pm 0.015$
B.1 Tagger

### v2

@@ 6 Conclusions
Departing from a state-of-the-art architecture for semantic dependency parsing, we were able to improve its performance on both semantic and syntactic dependency parsing on six datasets. On ERFGC, a dataset of directed acyclic semantic dependency graphs compiled from culinary recipes, our approach beats the state-of-the-art performance achieved by Bhatt et al. bhatt_end–end_2024 . Moreover, our results showed that a single BiLSTM layer can be sufficient to match or surpass the results of state-of-the-art architectures with a decrease in trained parameters of up to 85%. In the case of SciERC, a challenging dataset for semantic dependency parsing, we found that the performance boost was particularly great when training the parser with three BiLSTM layers. Moreover, for this challenging dataset, we find that scaling the predictions of the biaffine scorer can accelerate convergence speed even when fully fine-tuning models in the $100$ to $400M$ parameter range. In addition, for three of the datasets we also observed that the performance obtained with normalized and raw scores converged smoothly as the number of trained layers increased. This supported our claim that stacking BiLSTM layers mainly serves the purpose of producing an implicit regularization, and that this effect can be obtained by normalizing the scores, without any extra parameters. In additional experiments, we corroborated our findings using data in languages other than English nivre-etal-2020-universal , as well as non-linguistic data, for example using a dataset of molecule graphs ramakrishnan2014qm9 . Using GAT layers, we further confirmed our hypothesis that GNN-based iterative adjacency refinement ji_graph-based_2019 also benefits from normalization.
@@
In future work, we plan to extend our experiments to large-scale graph inference tasks which have been limited by the lack of parameter-efficient methods, such as long-form discourse parsing tasks. With regard to score scaling, we also wish to verify whether the positive effects of this normalization can carry over to other tasks, beyond latent graph inference. Finally, further explorations in model efficiency are warranted, given our findings, e.g., via pruning of model parameters.
Acknowledgements
We acknowledge the support of the Killam Foundation, the Natural Sciences and Engineering Research Council of Canada (NSERC), RGPIN-2022-03943, Canada Foundation of Innovation (CFI) and Research Nova Scotia. Advanced computing resources are provided by ACENET, the regional partner in Atlantic Canada, the Digital Research Alliance of Canada, and the Vector Institute.
Paolo Gajo carried out this work while visiting Dalhousie University and working with the Hypermatrix team. His work is funded by the EU (NextGenerationEU) with funds made available by the National Recovery and Resilience Plan (NRRP), Mission 4, Component 1, Investment 4.1 (M.D. 118/2023).
Appendix A Non-English data results
In this appendix, we carry out experiments on the SynDP task on six non-English datasets, all downloaded from Universal Dependencies. [^5] https://universaldependencies.org The datasets are listed in Table 4 and comprise the following languages: Arabic, Chinese, Italian, Japanese, Spanish, and Wolof. The results of the experiments are reported in Table 5. Since the data is in different languages, as encoder we use mBERT, the multilingual version of BERT. [^6] https://huggingface.co/google-bert/bert-base-multilingual-cased All model hyperparameters are kept the same as in the main setting.
Similarly to the English results observed for enEWT and SciDTB, the addition of BiLSTM layers increases the performance. Furthermore, additional layers have a diminished boosting effect on the performance. This shows that the benefits of normalization are independent of the language of which the dependencies are being parsed.
Table 6: Results for the non-linguistic datasets at varying depths of GAT layers.
$a$ $L_{\gamma}$ PCQM-Contact CIFAR10 Superpixel QM9
11 1 0.241 $\pm 0.044$ 0.407 $\pm 0.110$ 0.861 $\pm 0.008$
2 0.217 $\pm 0.019$ 0.528 $\pm 0.134$ 0.876 $\pm 0.006$
3 0.245 $\pm 0.034$ 0.751 $\pm 0.014$ 0.877 $\pm 0.008$
$\frac{1}{\sqrt{d}}$ 1 0.288 $\pm 0.090$ 0.531 $\pm 0.038$ 0.886 $\pm 0.001$
2 0.209 $\pm 0.028$ 0.585 $\pm 0.121$ 0.880 $\pm 0.004$
3 0.237 $\pm 0.019$ 0.703 $\pm 0.046$ 0.881 $\pm 0.013$
Appendix B Non-linguistic results
In this appendix we lay out the unlabeled latent graph inference experiments we carried out on three non-linguistic datasets: PCQM-Contact [10], CIFAR10 Superpixel [9], and QM9 [33]. PCQM-Contact and QM9 are molecule datasets, while CIFAR10 Superpixel is a dataset of superpixel clusters constructed from CIFAR10. Nodes in QM9 and PCQM-Contact respectively contain 9 and 11 features describing each atom in a molecule. In CIFAR10 Superpixel, nodes contain the 3D RGB features of each pixel cluster. In QM9 and CIFAR10 Superpixel, we concatenate the 3D and 2D spatial coordinates of the nodes to their features. Therefore, the resulting node feature dimensionalities of the three datasets are: $d_{\text{PCQM}}=9$, $d_{\text{CIFAR10}}=5$, and $d_{\text{QM9}}=14$.
We feed the graph node features into $L_{\gamma}\in\{1,2,3\}$ GAT layers and provide them with a fully-connected adjacency matrix without any edge attributes. Although the adjacency is fully connected, meaning each node can attend to all others in a single step, stacking $L_{\gamma}$ layers refines the globally aggregated representations over 1–3 iterations. The processed node features are passed into the biaffine layer to produce edge predictions. In this case, we pass the scores into a sigmoid function because we are not trying to predict an arborescence with the maximum spanning tree algorithm like in dependency parsing. We use $N=0$ BiLSTM layers to isolate the effect of the GAT layers with respect to any other network components. Scores are averaged over three different seeds.
The unlabeled F1 scores are reported in Table 6. In all cases, the results show that the gap in performance is large at $L_{\phi}=1$ and decreases (or flips) with additional layers. This corroborates the findings from the other settings and is an especially important result considering the phenomenon holds even for such a small network, consisting of just 100–160kk parameters.
Figure 5: Extended parser architecture, featuring $L_{\gamma}\in\{0,1,2,3\}$ pairs of biaffine and GAT layers.
Table 7: Micro-F1 (SemDP) and LAS (SynDP) for the labeled edge prediction task at varying depths of $L_{\gamma}\in\{0,1,2,3\}$ GAT layers. Best in bold.
$a$ $L_{\gamma}$ ADE CoNLL04 SciERC ERFGC enEWT SciDTB
11 0 0.517 $\pm 0.038$ 0.526 $\pm 0.046$ 0.123 $\pm 0.050$ 0.536 $\pm 0.013$ 0.610 $\pm 0.009$ 0.727 $\pm 0.006$
1 0.509 $\pm 0.022$ 0.493 $\pm 0.037$ 0.039 $\pm 0.020$ 0.599 $\pm 0.008$ 0.583 $\pm 0.011$ 0.731 $\pm 0.009$
2 0.509 $\pm 0.025$ 0.485 $\pm 0.033$ 0.029 $\pm 0.041$ 0.611 $\pm 0.015$ 0.540 $\pm 0.012$ 0.710 $\pm 0.013$
3 0.487 $\pm 0.040$ 0.481 $\pm 0.087$ 0.000 $\pm 0.000$ 0.585 $\pm 0.009$ 0.511 $\pm 0.015$ 0.700 $\pm 0.010$
$\frac{1}{\sqrt{d}}$ 0 0.556 $\pm 0.037$ 0.574 $\pm 0.028$ 0.156 $\pm 0.031$ 0.607 $\pm 0.009$ 0.671 $\pm 0.007$ 0.778 $\pm 0.006$
1 0.587 $\pm 0.015$ 0.550 $\pm 0.036$ 0.148 $\pm 0.028$ 0.639 $\pm 0.007$ 0.696 $\pm 0.005$ 0.809 $\pm 0.008$
2 0.534 $\pm 0.023$ 0.563 $\pm 0.029$ 0.128 $\pm 0.029$ 0.637 $\pm 0.007$ 0.657 $\pm 0.007$ 0.795 $\pm 0.008$
3 0.543 $\pm 0.024$ 0.514 $\pm 0.026$ 0.071 $\pm 0.021$ 0.616 $\pm 0.006$ 0.596 $\pm 0.010$ 0.770 $\pm 0.007$
Appendix C GAT results
In this appendix, we extend the main-setting parser, following [13], and carry out $k$-hop parsing experiments by passing the node representations through $L_{\gamma}\in\{0,1,2,3\}$ pairs of biaffine and GAT layers. These allow the model to encode higher-order dependencies before the final biaffine layer. Figure 5 depicts the extended architecture. We keep $N=0$ to isolate the effect of the GAT layers.
As Table 7 shows, the results of this architecture are consistent with the ones obtained using the main-setting 1-hop parser, with the performance increasing across the board when using normalization. As regards the number of GAT layers, performance increases when using $L_{\gamma}\in\{1,2\}$ layers, compared to $L_{\gamma}=0$, but drops again for all datasets at $L_{\gamma}=3$. In general, the best performance is obtained with normalization and $L_{\gamma}\in\{0,1\}$ layers. It is particularly interesting to notice that for enEWT the top performance without normalization is obtained at $L_{\gamma}=0$, while $L_{\gamma}=1$ provides better performance when normalizing biaffine scores. Overall, the usefulness of multi-hop dependency parsing for SemDP is scarce if compared with the SynDP results, where higher-order dependencies provide more useful information for the predictions.
Appendix D Ablations
This appendix provides a collection of ablation experiments for various components and settings of the model used in the main experiments.
@@ Appendix D Ablations
Table 8: Tagger ablations with best hyperparameters $(h_{\psi},d_{\text{MLP}})$. Best in bold, second-best underlined.
$a$ NN $\phi$ $\textbf{e}_{i}^{tag}$ ADE CoNLL04 SciERC ERFGC Mean
@@
11 0 $\blacksquare$ $\blacksquare$ 0.541 $\pm 0.021$ 0.399 $\pm 0.024$ 0.147 $\pm 0.049$ 0.548 $\pm 0.010$ 0.515
1 $\blacksquare$ $\blacksquare$ 0.657 $\pm 0.011$ 0.556 $\pm 0.021$ 0.282 $\pm 0.009$ 0.676 $\pm 0.010$
2 $\blacksquare$ $\blacksquare$ 0.667 $\pm 0.011$ 0.573 $\pm 0.025$ 0.273 $\pm 0.010$ 0.694 $\pm 0.010$
3 $\blacksquare$ $\blacksquare$ 0.662 $\pm 0.027$ 0.562 $\pm 0.021$ 0.299 $\pm 0.023$ 0.705 $\pm 0.011$
$\frac{1}{\sqrt{d}}$ 0 $\blacksquare$ $\blacksquare$ 0.567 $\pm 0.014$ 0.438 $\pm 0.033$ 0.181 $\pm 0.027$ 0.612 $\pm 0.008$ 0.541
1 $\blacksquare$ $\blacksquare$ 0.668 $\pm 0.017$ 0.597 $\pm 0.015$ 0.299 $\pm 0.019$ 0.692 $\pm 0.009$
2 $\blacksquare$ $\blacksquare$ 0.676 $\pm 0.019$ 0.596 $\pm 0.014$ 0.312 $\pm 0.011$ 0.699 $\pm 0.009$
3 $\blacksquare$ $\blacksquare$ 0.686 $\pm 0.025$ 0.602 $\pm 0.017$ 0.320 $\pm 0.013$ 0.708 $\pm 0.008$
11 0 $\blacksquare$ $\square$ 0.543 $\pm 0.013$ 0.402 $\pm 0.023$ 0.162 $\pm 0.019$ 0.554 $\pm 0.008$ 0.517
1 $\blacksquare$ $\square$ 0.674 $\pm 0.011$ 0.527 $\pm 0.026$ 0.275 $\pm 0.013$ 0.677 $\pm 0.005$
2 $\blacksquare$ $\square$ 0.672 $\pm 0.019$ 0.575 $\pm 0.009$ 0.293 $\pm 0.012$ 0.689 $\pm 0.011$
3 $\blacksquare$ $\square$ 0.657 $\pm 0.027$ 0.583 $\pm 0.011$ 0.301 $\pm 0.012$ 0.697 $\pm 0.007$
$\frac{1}{\sqrt{d}}$ 0 $\blacksquare$ $\square$ 0.563 $\pm 0.013$ 0.443 $\pm 0.019$ 0.188 $\pm 0.006$ 0.612 $\pm 0.003$ 0.534
1 $\blacksquare$ $\square$ 0.653 $\pm 0.023$ 0.565 $\pm 0.027$ 0.299 $\pm 0.020$ 0.687 $\pm 0.006$
2 $\blacksquare$ $\square$ 0.672 $\pm 0.017$ 0.592 $\pm 0.020$ 0.307 $\pm 0.008$ 0.702 $\pm 0.005$
3 $\blacksquare$ $\square$ 0.661 $\pm 0.022$ 0.593 $\pm 0.017$ 0.305 $\pm 0.014$ 0.708 $\pm 0.007$
11 0 $\square$ $\blacksquare$ 0.550 $\pm 0.026$ 0.387 $\pm 0.030$ 0.157 $\pm 0.012$ 0.553 $\pm 0.006$ 0.514
1 $\square$ $\blacksquare$ 0.667 $\pm 0.022$ 0.532 $\pm 0.036$ 0.273 $\pm 0.013$ 0.673 $\pm 0.006$
2 $\square$ $\blacksquare$ 0.665 $\pm 0.021$ 0.565 $\pm 0.024$ 0.278 $\pm 0.027$ 0.694 $\pm 0.013$
3 $\square$ $\blacksquare$ 0.676 $\pm 0.021$ 0.558 $\pm 0.051$ 0.288 $\pm 0.012$ 0.706 $\pm 0.006$
$\frac{1}{\sqrt{d}}$ 0 $\square$ $\blacksquare$ 0.578 $\pm 0.022$ 0.437 $\pm 0.030$ 0.187 $\pm 0.014$ 0.611 $\pm 0.008$ 0.534
1 $\square$ $\blacksquare$ 0.651 $\pm 0.032$ 0.559 $\pm 0.029$ 0.301 $\pm 0.007$ 0.684 $\pm 0.003$
2 $\square$ $\blacksquare$ 0.673 $\pm 0.029$ 0.582 $\pm 0.017$ 0.310 $\pm 0.014$ 0.701 $\pm 0.008$
3 $\square$ $\blacksquare$ 0.659 $\pm 0.019$ 0.586 $\pm 0.026$ 0.324 $\pm 0.019$ 0.708 $\pm 0.012$
11 0 $\square$ $\square$ 0.547 $\pm 0.019$ 0.402 $\pm 0.033$ 0.156 $\pm 0.029$ 0.549 $\pm 0.015$ 0.516
1 $\square$ $\square$ 0.651 $\pm 0.017$ 0.552 $\pm 0.012$ 0.288 $\pm 0.008$ 0.680 $\pm 0.007$
2 $\square$ $\square$ 0.664 $\pm 0.031$ 0.566 $\pm 0.012$ 0.288 $\pm 0.017$ 0.693 $\pm 0.011$
3 $\square$ $\square$ 0.663 $\pm 0.008$ 0.561 $\pm 0.014$ 0.291 $\pm 0.012$ 0.703 $\pm 0.007$
$\frac{1}{\sqrt{d}}$ 0 $\square$ $\square$ 0.566 $\pm 0.014$ 0.431 $\pm 0.015$ 0.182 $\pm 0.023$ 0.606 $\pm 0.012$ 0.534
1 $\square$ $\square$ 0.655 $\pm 0.015$ 0.567 $\pm 0.010$ 0.286 $\pm 0.018$ 0.678 $\pm 0.013$
2 $\square$ $\square$ 0.678 $\pm 0.014$ 0.591 $\pm 0.026$ 0.307 $\pm 0.007$ 0.702 $\pm 0.007$
3 $\square$ $\square$ 0.684 $\pm 0.022$ 0.595 $\pm 0.017$ 0.321 $\pm 0.016$ 0.698 $\pm 0.015$
D.1 Tagger